OpenStudy (imamermaid29):

To use a vacuum cleaner at a car wash, customers must pay an initial $3.00 and $0.25 for each 15 min interval. Which table shows how much it costs to use the vacuum cleaner for different amounts of time? A. Vacuum Cleaner Costs 15 min intervals 3 5 7 9 Cost $0.75 $1.25 $1.75 $2.25 B. Vacuum Cleaner Costs 15 min intervals 3 5 7 9 Cost $3.75 $4.25 $4.75 $5.25 C. Vacuum Cleaner Costs 15 min intervals 3 5 7 9 Cost $14.25 $20.50 $29.25 $36.75 D. Vacuum Cleaner Costs 15 min intervals 3 5 7 9 Cost $15.75 $16.25 $16.75 $17.2

OpenStudy (imamermaid29):

@ooxBRITTANYxoo

OpenStudy (imamermaid29):

@cobra-strikes

OpenStudy (imamermaid29):

@green_1

OpenStudy (texaschic101):

I think B

OpenStudy (green_1):

it is b i just got the answer

OpenStudy (word2):

(x) = 0.25x + 3 x = number of 15 minute intervals f(3) = 0.25(3) + 3 f(3) = 0.75 + 3 f(3) = 3.75 f(5) = 0.25(5) + 3 f(5) = 1.25 + 3 f(5) = 4.25 f(7) = 0.25(7) + 3 f(7) = 1.75 + 3 f(7) = 4.75 f(9) = 0.25(9) + 3 f(9) = 2.25 + 3 f(9) = 5.25 the answer is B

OpenStudy (green_1):

any other questions